# Enhancing Efficiency in Multidevice Federated Learning through Data Selection

### v4

@@ Abstract.
Federated learning (FL) in multidevice environments creates new opportunities to learn from a vast and diverse amount of private data. Although personal devices capture valuable data, their memory, computing, connectivity, and battery resources are often limited. Since deep neural networks (DNNs) are the typical machine learning models employed in FL, there are demands for integrating ubiquitous constrained devices into the training process of DNNs. In this paper, we develop an FL framework to incorporate on-device data selection on such constrained devices, which allows partition-based training of a DNN through collaboration between constrained devices and resourceful devices of the same client. Evaluations on five benchmark DNNs and six benchmark datasets across different modalities show that, on average, our framework achieves $\sim$19higher accuracy and $\sim$58lower latency; compared to the baseline FL without our implemented strategies. We demonstrate the effectiveness of our FL framework when dealing with imbalanced data, client participation heterogeneity, and various mobility patterns. As a benchmark for the community, our code is available at github.com/dr-bell/data-centric-federated-learning.
copyright: none
@@ 1. Introduction
Multidevice FL enables learning new insights from various clients by training DNNs on personal data in a more private manner (McMahan et al., 2017a; Kairouz et al., 2021). However, some wearable or embedded devices have limited computing resources, little on-device storage, and variable network connectivity. Most existing FL frameworks do not account for constraints of personal devices and instead assume either modern smartphones (Hard et al., 2018; Niu et al., 2020) or accelerators (Cho et al., 2022; Zhang et al., 2022) as target client devices. Although smartphones and accelerators are less powerful than cloud machines, they are nevertheless equipped with mobile GPUs, possess gigabytes of runtime memory, and have fairly stable connectivity to a central server, which simplifies the requirements of DNN training (Yang et al., 2021; Lin et al., 2020).
Challenge. We focus on multidevice FL for ubiquitous constrained devices (UCDs). Examples are wearable devices such as earbuds, glasses, or rings, and embedded devices such as environmental or industrial sensors. We are driven by the motivation that UCDs are becoming the primary data-producing devices for individuals and industries due to various on-device sensors, such as cameras and inertial measurement units. UCDs offer a wealth of spatio-temporal data that is frequently absent when focusing solely on smartphones or plugged-in devices. For instance, people often opt for wristbands, smart rings, or earbuds over carrying a smartphone in their hand or pocket while engaging in outdoor, industrial, or sports activities (Seneviratne et al., 2017; Mo and Zhou, 2021). Moreover, low-cost sensors distributed across farms or forests remotely monitor crops and livestock and detect potential threats (Papst et al., 2021). Such abundance and diversity in sensory data captured by UCDs can facilitate the learning of more effective models for emerging applications in personal and industrial environments. Nevertheless, there are two challenges for FL on UCDs:
The first challenge is the network connectivity to a central server. Usually, UCDs use an access point (AP) over wireless channels, e.g., a smartphone or a router, as a relay to connect to the central server. Such wireless connections, e.g., via Bluetooth Low Energy (BLE), are unreliable due to the devices’ mobility. Particularly for outdoor settings, UCDs may not always be connected to the companion AP, and hence therefore no immediate connection to the central server.
The second challenge emerges from constraints on the memory and compute resources available on UCDs, particularly the model size and functionality of DNNs that can be trained locally on UCDs. Although a naive solution is to transfer all data from UCDs to their respective APs and perform training of DNNs on APs, storing all data could exceed the storage capacity of UCDs, especially when UCDs disconnect from their AP often. To tackle these challenges effectively, we require FL frameworks that dynamically distribute data and computations between UCDs and APs.
Solution. We propose an end-to-end FL framework that orchestrates the local training of DNNs among UCDs and APs by integrating on-device data selection alongside partition-based model training. We refer to our proposed framework as Centaur. (A) Considering computation constraints, Centaur initializes the DNN in two partitions: an encoder (i.e., feature extractor) to be only trained on APs, followed by a lightweight classifier to be trained on both APs and UCDs (§3.2).
This is inspired by split (Gupta and Raskar, 2018; Thapa et al., 2022) and transfer (Weiss et al., 2016) learning, where UCDs keep the encoder frozen and only train the classifier, while the AP train the full DNN. (B) Considering memory and connectivity constraints, Centaur performs data selection (Jiang et al., 2019) by analyzing the training loss and the gradients norm of the classifier part, to decide which data points captured by UCDs contribute more to the training of which part of the DNN. Here, data points are categorized as either of (i) discarded if they have very low loss values, (ii) kept locally on the UCD to train the classifier part if their loss values or gradients norm are not very high, and (iii) transmitted to the AP to train both encoder and classifier part if they cause high values for both loss and gradients norm [^1] The assumption is each pair of UCD and AP belongs to the same client, thus transmitting data from a UCD to its AP does not violate the clients’ privacy. (§3.3).
Centaur differs from the prior works on hierarchical FL (Liu et al., 2020; Abad et al., 2020), since they neither perform partition-based training nor incorporate data selection. This makes our FL framework adaptive to the available resources in a multidevice setting.
Evaluation. We perform evaluations on five benchmark DNNs and six benchmark datasets. We show that compared to existing FL alternatives, Centaur not only saves more bandwidth by reducing the communication cost associated with offloading samples from the UCD to the AP, but also reduces the latency of training, and as a result, the energy consumption, on both the AP and the UCD. As a prime example of our experimental results, Figure 1 shows the latency vs. accuracy of performing FL for two different tasks: (1) four benchmark DNNs, with their encoder pre-trained on ImageNet (Deng et al., 2009) , and then finally trained for CIFAR10 (Krizhevsky et al., 2009) image classification with 100 clients, and (2) a benchmark ConvNet (Chang et al., 2020), trained from scratch, on three human-activity recognition (HAR) datasets with 20 clients. In both figures, we compare Centaur with the baseline of standard FL on UCDs without our implemented strategies. Centaur, through data selection and partition-based model training, achieves up to $19\%$ higher accuracy and $58\%$ lower latency, on average. We provide thorough experiments and comparisons with other alternative baselines in §6.
@@
Figure 1. Training latency vs. test accuracy of Centaur compared with standard UCD training (i.e., only training the classifier on UCDs without data selection and partition-based training). (Left) Four benchmark pre-trained DNNs fine-tuned on CIFAR10 dataset with FL. (Right) A ConvNet trained with FL on three benchmark wearable sensor datasets: MotionSense, PAMAP2, and UCIHAR.
Contribution. Our solution enables UCDs to efficiently participate in multidevice FL alongside their upstream APs, by incorporating a customizable data selection scheme and a partition-based training regime. The impact is a decrease in computational and communication costs, as well as an improvement in the model’s test accuracy. Our experiments on a small testbed of RaspberryPis validate that Centaur is promising in decreasing overall storage requirements and training time due to employing effective data selection. Moreover, our analysis on the mobility of UCDs accounts for the real-world spatiotemporal mobility of edge devices participating in FL. Our extensive empirical analyses, including measurement of cost, imbalance data, participation heterogeneity, and connection probability, indicate that Centaur achieves higher efficiency (i.e., high accuracy and low cost) across most scenarios. Our code is open-source to encourage further advancements in this domain.
2. Identifying Design Challenges: A Motivational Study
Terminology. A ubiquitous constrained device (UCD) is a mobile device with restricted hardware and connectivity resources. An access point (AP) is a resourceful device that has much better hardware and connectivity capabilities than its corresponding UCD. A server is a central entity owned by a service provider that organizes FL over several clients. An encoder ($E$) is the first part of the DNN and extracts the features of input data. A classifier ($C$) is the second part of the DNN and classifies the encoded features into a final prediction. Let $\mathbf{X}$ denote the input data and $\mathbf{y}$ denote its label.
2.1. The Study
We perform measurement experiments to identify the key design requirements that can facilitate the deployment of FL on constrained devices.
@@ 2.1. The Study
Setup. We create a testbed with four RaspberryPi 4 Model B as the clients (Velasco-Montero et al., 2018; Drakopoulos et al., 2019); all running the 64-bit Raspbian OS with total secondary storage of $32$GB. We set up a cohort of clients having varying primary memories, with clients C1, C2, C3, and C4 having $1$GB, $2$GB, $4$GB, and $2$GB of primary memory, respectively. For analyzing the CPU and memory consumption of the clients, we use the logs from the /proc/stat and free system calls, respectively. For the model architecture, we use a benchmark architecture MobileNetV3 (Howard et al., 2019) (pre-trained on ImageNet) as the encoder followed by a medium-sized classifier with two fully connected layers with an intermediate drop-out layer. We use CIFAR10 dataset and measure the resource and time consumed on-device for both training, per each epoch, and inference. Finally, for implementing FL setup, we use Flower (Beutel et al., 2022) over four RaspberryPi devices and a Mac Mini as the server, as shown in Figure 2a.
@@
Running Standard FL. We implement standard FL (for $100$ rounds) over four RaspberryPi clients with the standard FedAvg (McMahan et al., 2017b). We consider two scenarios: (1) when only the classifier part is trained, and (2) when the entire model (encoder and classifier) is trained. For the first scenario, we observed in our experiments that the standard FL runs successfully over all four clients. However, for the second scenario, client C1 with 1GB of primary memory fails to participate in the FL process. Notably, 1GB of primary memory is still significant in terms of memory capability and typical UCDs usually have a much lesser amount of memory than this. Therefore, it is clear that in such cases, it is much more challenging to deploy standard FL over UCDs. To further assess this, we conduct a series of experiments in the following section to investigate the resource and time consumption for training DNNs on our testbed.
On-device Training. Considering the challenges discussed previously, we analyze the latency and resource consumption for local training of MobileNetV3. Figure 2 shows that there is a significant rise in the overall CPU and memory consumption when the entire model (encoder and classifier) is trained on-device in comparison to when only the classifier is trained and the encoder kept frozen. A key observation is that by training only the classifier, we require less processing power and approximately the same amount of memory resources as when the entire model is used solely for inference. Moreover, Figure 3 shows that the time taken to train the classifier is less than the time to perform a single forward pass on the entire model. These measurements confirm that training the classifier only is 50×50$\times$50 × time efficient compared to training the entire model, which helps in reducing the battery consumption significantly. When we follow a naive approach of randomly dropping samples from the training dataset, we observe a decline in the overall training time of the classifier (see Figure LABEL:fig:select_train). However, this comes at the cost of reduced accuracy, which might be detrimental to the overall system.
Figure 2. Resource consumption of training DNNs on (a) the designed setup over four RaspberryPi 4 Model B. We compare (b) CPU Consumption and (c) Memory Consumption for three scenarios when: (i) only the classifier is updated, (ii) the whole model is updated, and (ii) the whole model is used only at inference mode.
2.2. Insights
Here, we summarize the understandings derived from our motivational study:
(1) Resource and Training Efficiency. As UCDs typically have limited memory, training on the device should be memory efficient. This motivates using a frozen pre-trained encoder, and only training the classifier part on the device (Mathur et al., 2021). However, freezing the weights of the encoder limits the expressiveness of the model and prevents the model from learning important features of the newly collected data over the UCDs. Thus, while minimizing the on-device model training on the UCDs, our system should ensure that the encoder is periodically trained on a portion of new data collected from the UCDs, potentially on a more resourceful device where the entire model can be trained.
(2) Data Efficiency and Spatiotemporal Coverage. Typically, UCDs are deployed to maximize the spatiotemporal coverage of activities and events. Hence, our design should leverage the spatiotemporal richness of such data to train a better model. Such spatiotemporal diversity also makes UCDs more prone to intermittent connectivity, which may require the framework to store training data on-device. UCDs possess limited secondary storage, and even when they have persistent connectivity to the AP, streaming high volumes of data could drain their battery. Thus, our design should dynamically select a small subset of data points for storing locally or transmitting to the APs. This needs to be designed in a sophisticated manner such that it does not compromise the overall performance of the system.
Figure 3. (a) Time for backpropagation across (i) the entire model (encoder and classifier), (ii) across the classifier only, and (iii) for forward pass on the entire model. (b) The trend of training time and accuracy (for one epoch) for different sizes of training sets.
3. Our Proposed Framework: Centaur
With our understanding developed from the motivational study, we introduce our proposed framework. Figure 4 gives a schematic overview, and below we elaborate on the algorithms used in Centaur.
3.1. The Overview
The first level includes UCDs, which are connected to their APs via BLE, Wi-Fi, Backscatter, or other wireless technologies. We assume the connection between UCD and AP is not always available or reliable due to the natural mobility of UCDs. We consider that the data selection procedure and the training of the classifier only happen on UCDs. The second level, which includes APs, is responsible for training the encoder in conjunction with the classifier atop it. The third level includes a central server, which performs model aggregation for both classifiers and encoders that are trained by UCDs and APs. Here, we present a step-by-step description of our proposed framework.
Step \⃝raisebox{-0.9pt}{1}: We assume that the server receives the information about UCDs’ resource. Accordingly, the server initializes a DNN model $M$ that is formed of an encoder $E$ and a classifier $C$, and then distributes the $E$ and $C$ to the APs and subsequently to UCDs. We explain model initialization details in §3.2.
Step \⃝raisebox{-0.9pt}{2}: UCDs start FL by performing data selection on the collected data $D$, based on the computed loss value via the forward pass and the magnitude of the gradient of the classifier’s last layer. Each sample point will be categorized in one of (i) $D_{N}$ to be discarded, (ii) $D_{C}$ for locally training $C$ on the UCD, or (iii) $D_{M}$ to be transmitted to AP for training $M$. We explain data selection details in §3.3.
Step \⃝raisebox{-0.9pt}{3}: The UCD discards data $D_{N}$ in the current epoch without any further training steps or backward pass on them.
Step \⃝raisebox{-0.9pt}{4}: The UCD performs backward pass through the classifier $C$ on data $D_{C}$, while the encoder $E$ is frozen. Then, the UCD updates the parameters of $C$ via computed gradients. Notice that, in practice, the gradient computation done in Step \⃝raisebox{-0.9pt}{2} can be reused for this update.
Step \⃝raisebox{-0.9pt}{5}: UCDs share their updated classifier $C$ to the server, and the server performs aggregation on all $C$ from participating UCDs to obtain one aggregated classifier.
Step \⃝raisebox{-0.9pt}{6}: The server sends aggregated classifier back to the UCDs and APs.
Step \⃝raisebox{-0.9pt}{7}: With the updated classifier, the AP trains the encoder $E$ together with the classifier $C$ atop $E$ on data $D_{M}$. That is, $D_{M}$ is fed for a forward pass and then a backward pass throughout to update $M$.
Step \⃝raisebox{-0.9pt}{8}: The updated $M$ (including $E$ and $C$) is shared from the AP to the server. After collecting updated $M$ from all participating APs, the server performs model aggregation to obtain the aggregated encoder $E$ and classifier $C$ and distribute them back to UCDs and APs as in Step \⃝raisebox{-0.9pt}{6}. See § 3.4 for more details of FL training procedures. Steps from \⃝raisebox{-0.9pt}{2} to \⃝raisebox{-0.9pt}{8} will be repeated for several FL rounds to update $C$ on UCDs and $E$ on APs.
Figure 4. The overview of our framework, including model initialization (§ 3.2), data selection (§ 3.3), and partition-based training and aggregation (§ 3.4). We explain the set up in § 3.1 and elaborate the details of Steps \⃝raisebox{-0.9pt}{1} to \⃝raisebox{-0.9pt}{8}.

### v6

@@ Abstract.
UbiquitousTo be presented in the 10th ACM/IEEE Symposium on Edge Computing (SEC’25), December 2025, Arlington, VA, USA. wearable and mobile devices provide access to a diverse set of data. However, the mobility demand for our devices naturally imposes constraints on their computational and communication capabilities. A solution is to locally learn knowledge from data captured by ubiquitous devices, rather than to store and transmit the data in its original form. In this paper, we develop a federated learning framework, called Centaur, to incorporate on-device data selection at the edge, which allows partition-based training of a deep neural nets through collaboration between constrained and resourceful devices within the multidevice ecosystem of the same user. We benchmark on five neural net architecture and six datasets that include image data and wearable sensor time series. On average, Centaur achieves $\sim$19% higher classification accuracy and $\sim$58% lower federated training latency, compared to the baseline. We also evaluate Centaur when dealing with imbalanced non-iid data, client participation heterogeneity, and different mobility patterns. To encourage further research in this area, we release our code at github.com/nokia-bell-labs/data-centric-federated-learning.
Key words and phrases:
Federated Learning, Constrained Devices, Data Selection, Partition-Based Training, Edge Intelligence, On-device AI
@@ 1. Introduction
Figure 1. The latency of federated training versus classification accuracy on test dataset. We compare Centaur with standard federated training that only trains the classifier on ubiquitous constrained devices (UCDs) without data selection and partition-based training. (Top plot) Four models were tested on image data from CIFAR-10, and (bottom plot) a model was tested on three datasets of wearable sensor time-series.
With the growing trend of utilizing ubiquitous devices in personal and industrial environments, federated learning (FL) enables the discovery of new patterns in personal data by training deep neural networks (DNNs) on data captured by various users in a more private manner floriginal; flsurvey. Especially, in a multidevice ecosystem, the owner can seamlessly share data across the devices within their trusted environment while protecting data privacy cho2022flame. This allows us to utilize the capabilities of multiple devices to participate in the FL process. However, wearable and ubiquitous devices have limited computing resources, little on-device storage, and inconsistent network connectivity.
Most of the existing FL frameworks do not account for constraints of personal devices and instead assume either modern smartphones hard2018federated; niu2020billion, or edge devices with DNN accelerators cho2022flame; zhang2022federated, as target client devices. Although smartphones and accelerators are less powerful than cloud machines, they are nevertheless equipped with mobile GPUs, possess gigabytes of runtime memory, and have fairly stable connectivity to a central server, which simplifies the requirements of DNN training yang2021flop; lin2020ensemble.
In this paper, we propose a novel FL framework to include ubiquitous constrained devices (UCDs), i.e., wearable devices such as earbuds, glasses, or rings, or distributed edge devices such as environmental cameras or industrial sensors. We believe UCDs are becoming the primary data-producing devices for both individuals and industries, due to various on-device sensors, such as cameras or inertial measurement units. These devices offer a wealth of spatiotemporal data that is frequently absent when focusing solely on smartphones or plugged-in devices. For example, people often opt for wristbands, smart rings, or earbuds over carrying a smartphone in their hand or pocket while engaging in outdoor, industrial, or sports activities seneviratne2017survey; mo2021adapting, and sensors distributed across farms or forests remotely monitor crops and livestock to detect hazards papst2021exploring. The main motivation of our work is the fact that the abundance and diversity in data captured by ubiquitous devices can facilitate learning more effective models for emerging applications in personal and industrial environments.
There are several significant challenges. Usually, distributed or wearable UCDs use an access point (AP), like a smartphone or a router, as a relay to connect to a central server in the cloud. Data captured by a UCD is communicated to the AP using communication protocols like Bluetooth Low Energy (BLE). In outdoor settings under mobility, UCDs may not always be connected to the companion AP, thus inaccessible to the central server. Beside this, the constrained data storage and computing resources will restrict the size and functionality of DNNs that can be trained locally on UCDs. A naive solution is to transfer all data from UCDs to their respective APs and perform training of DNNs on APs. Yet, storing all data could exceed the device’s memory, especially when UCDs disconnect from their AP. To demonstrate these points, in §3, we conduct a motivating study on a real testbed to identify the essential design requirements that can enable the deployment of FL on resource-constrained devices. To tackle these challenges effectively, we believe FL solutions relying on on-device training should dynamically distribute data and computations between UCDs and their companion APs.
Table 1. Comparison with related work. (1) Model Partitioning splits the model into an encoder and a classifier. (2) Data Selection trains on a subset of available data. (3) On-device Training considers backpropagation on constrained devices. (4) Spatiotemporal Coverage includes mobility patterns of portable devices.
Model Partitioning Data Selection On-device Training Spatiotemporal Coverage
Multitier FL abad2020hierarchical; liu2020client; abdellatif2022communication; cho2022flame; gupta2018distributed × × × ×
Split FL diao2020heterofl; zhang2021fedzkt; zhang2021parameterized; li2021hermes; dai2022dispfl; bibikar2022federated × × ×
On-device Training lee2019intermittent; islam2019zygarde; fedorov2019sparse; banbury2021micronets; xu2022etinynet; liberis2021munas; lin2022device; kopparapu2022tinyfedtl × ×
Centaur (ours)
In §4 and Figure 3, we introduce Centaur: a federated learning framework that orchestrates local training among UCDs and APs by integrating on-device data selection with partition-based model training.
(A) Considering computation constraints, we initialize the DNN in two partitions: an encoder (i.e., feature extractor) to be only trained on APs, followed by a lightweight classifier to be trained on both APs and UCDs (§4.1). This allows Centaur to utilize all the data available on UCDs for fine-tuning the DNN while occasionally storing and transmitting a portion of the data to the APs to re-train the entire DNN and adjust to the runtime variability in data distributions gupta2018distributed; thapa2022splitfed; weiss2016survey.
(B) Considering memory and connectivity constraints, we perform data selection jiang2019accelerating by analyzing the training loss and the gradients norm of the classifier part, to decide which data points captured by UCDs contribute more to the training of which part of the DNN. Through data selection, data points are categorized as either of (i) discarded if they have very low loss values, (ii) kept locally on the UCD to train the classifier part if their loss values or gradients norm are not very high, and (iii) transmitted to the AP to train both encoder and classifier part if they cause high values for both loss and gradients norm (§4.2). We assume that each pair of UCD and AP belongs to the same client, thus transmitting data from a UCD to its AP does not violate the clients’ privacy.
(C) We benchmark on five DNNs architectures and six datasets of two different modalities, and compared to existing FL alternatives. Our results show that Centaur saves more bandwidth by reducing the communication cost associated with offloading samples from a UCD to the AP, and also reduces the latency of training (and accordingly the energy consumption) on both UCD and AP. As a prime example of our experimental results, Figure 1 shows the training latency versus classification accuracy of performing FL for two different tasks: (1) four benchmark DNNs, with their encoder pre-trained on ImageNet deng2009imagenet, and then trained for CIFAR10 krizhevsky2009learning image classification with 100 FL clients, and (2) a benchmark ConvNet chang2020systematic, trained from scratch, on three human-activity recognition datasets with 20 FL clients. In both figures, we compare Centaur with the baseline of standard FL on UCDs without our implemented strategies. Centaur, through data selection and partition-based model training, achieves up to $19\%$ higher accuracy and $58\%$ lower latency, on average.
The key contribution of our work is a better integration of ubiquitous constrained devices into federated learning by leveraging the advantages of multi-device ecosystems. We achieve this by introducing a customizable data selection scheme and a partition-based training approach, which collectively reduce computational and communication costs while enhancing model accuracy. Our experiments demonstrate that Centaur effectively minimizes storage requirements and training time through efficient data selection. Furthermore, our analysis accounts for the real-world spatiotemporal mobility of edge devices in FL. Empirical evaluations, covering cost, data imbalance, participation heterogeneity, and connection probability, show that Centaur consistently delivers higher efficiency, achieving improved accuracy at lower costs across diverse scenarios.
2. Related Work
Table 1 compares Centaur with major prior work on different aspects of multidevice FL. Compared to the standard edge–server FL, edge–access point–server FL liu2020client; abad2020hierarchical considers a middle layer, e.g., a cellular base station, to orchestrate FL training across clients and a central server. Clients can participate with several devices cho2022flame; diao2020heterofl; zhang2021fedzkt, which achieve better trade-offs between accuracy and bandwidth consumption. However, these works assume the local training of the entire model on edge devices, without any resource limitations. They also do not utilize the potential of resource-rich devices to assist a resource-constrained device during training. They use APs only for aggregation and communication, which makes FL impractical in scenarios involving UCDs.
Other works consider inference only (i.e., forward pass) fedorov2019sparse; banbury2021micronets; xu2022etinynet; liberis2021munas, sparse training li2021hermes; bibikar2022federated, or heterogenous architectures across clients of different capabilities kang2023nefl; park2024federated. Sparse training or pruning requires computational resources (especially memory) that UCDs do not have. Few works propose solutions for full on-device training (i.e., forward and backward pass) lin2022device, but not for FL scenarios as they often make impractical assumptions like the availability of large datasets for network architecture search liberis2021munas, or unconstrained memory for training the full model diao2020heterofl; zhang2021fedzkt. Transfer learning can be used on UCDs to fix the encoder architecture, which in turn restricts the extraction of newer features, and only trains the last layer for the personalized training zhang2021fedzkt.
Without utilizing the power of APs, training the entire model on UCDs requires a large sample size to learn generalized features li2021hermes; bibikar2022federated. Simply considering different sub-models diao2020heterofl; kang2023nefl; park2024federated requires more sophisticated aggregation strategies and shallower models for resource-constrained devices, which may not be capable of learning the complex patterns from the data. Recent works enable training on micro-controllers with limited memory, such as TensorFlow Lite Micro david2021tensorflow or Tiny Training Engine lin2022device. These works do not consider FL as a potential use case and do not account for the advantages of a multidevice ecosystem.
3. A Motivating Study
As depicted in Figure 3, a ubiquitous constrained device (UCD) is a small, low-power portable device with restricted hardware and connectivity resources. Examples include smart rings, earbuds or glasses (personal wearables) and tiny environmental or industrial sensors. Each UCD connects to the user’s personal access point (AP), usually over a low-power link such as Bluetooth Low Energy. An AP is a resourceful device that has much better hardware and connectivity capabilities than its corresponding UCD (for instance, a smartphone, home router, or gateway). We assume the AP is always part of the user’s device ecosystem that allows the UCD to transmit data to the AP without violating the user’s privacy. We generally assume one AP serves multiple UCDs (e.g., a person’s phone might pair with a smartwatch, earbuds, and other sensors simultaneously). Indeed, modern homes routinely host on the order of 20–30 connected devices, so a single AP often handles many UCDs.
A server is a central entity owned by a service provider that orchestrates FL over several clients. Considering a deep neural network architecture, an encoder ($E$) is the first part of the DNN that extracts the features from input data. A classifier ($C$) is the second part of the DNN, which performs the final classification using the encoded features. We consider supervised learning tasks where $\mathbf{X}$ denote the input data and $\mathbf{y}$ denote its label.
@@ 3. A Motivating Study
Setup. To identify the key design choices, we create a testbed with four RaspberryPi 4 Model B as edge devices velasco2018performance; drakopoulos2019real; all running the 64-bit Raspbian OS with total secondary storage of 3232GB. We set up four clients with varying primary memories: C1, C2, C3, and C4 having 11GB, 22GB, 44GB, and 22GB, respectively. Indeed, Raspberry Pi devices are more capable than many wearables; however, they serve as a widely adopted, reproducible proxy for constrained clients cho2022flame, enabling controlled CPU/memory benchmarking while we contextualize our findings with real-world wearable hardware specifications and multitasking considerations. For analyzing the CPU and memory consumption of the clients, we use logs of /proc/stat and free operating-system calls, respectively. We use MobileNetV3 howard2019searching, pre-trained on ImageNet, as the encoder followed by two fully-connected layers with an intermediate drop-out layer. Using CIFAR10, we measure the resources and time needed for both on-device training (per epoch) and inference. We implement FL with Flower beutel2022flower and use a desktop computer as the server, as shown in Figure 2a.
Figure 2. On four RaspberryPis, we measure memory consumption of running FL for when (i) only the classifier is trained, (ii) the entire model is trained, and (iii) the entire model runs inference only.
[…]
@@
Key Observations. We run FedAvg floriginal for $100$ rounds with four RaspberryPi devices, considering two scenarios: (1) when only the classifier part is trained, and (2) when the entire model (encoder and classifier) is trained. When running our experiments for the first scenario, we observed that the standard FL runs successfully over all four clients. However, for the second scenario, client C1 with 1GB of primary memory always failed to participate in the FL process. This is primarily due to the significant rise in the CPU and memory consumption when the entire model (both encoder and classifier) are trained on the device (see Figure 2).
Notably, 1GB of primary memory is still significant in terms of memory capability, and many UCDs may have significantly less memory than this. Therefore, it is clear that in such cases, it is much more challenging to deploy standard FL over UCDs. Moreover, training only the classifier consumes far less resources and is even comparable to the resource usage required for a forward pass. We noticed a similar pattern for time consumption, where training the classifier only is $50\times$ time efficient compared to training the entire model. One alternative solution to decrease the overall training time is to reduce the training data size by randomly discarding samples. However, we observe that such an approach comes at a cost of significantly reduced accuracy, which is detrimental to the overall system.
4. Method
In Figure 3, we introduce Centaur, our federated learning framework. Here, we elaborate on each component in detail.
[…]
Since full backpropagation involves a forward pass, the storage of all intermediate activations, and a backward pass through every layer to compute gradients, it is the most compute- and memory-intensive phase of model training. Centaur mitigates this by freezing the deep encoder and training only the lightweight classifier head on UCDs. Each device performs a single forward pass through the fixed encoder and backpropagates only through the shallow fully-connected classifier, which has orders of magnitude fewer parameters. This design significantly reduces the on-device workload. Our motivationa study in §3 confirmed this trend, with classifier-only training running several times faster and using far less memory than updating the entire network. Thus, Centaur enables effective training on local data while keeping UCD training lightweight and efficient.